Calculate the pH for each of the cases in the titration of 25 mL of .120 M pyridine, with .123 M HBr.

C. after addition of 20.0mL of HBr 5.33

D. after addition of.25.0 mL of HBr

E. after addition of 32.0 mL of HBr
Answer #1

pKb of pyridine = 8.77

pKa = 5.23

C. after addition of 20.0mL of HBr

mmoles of base pyridine = 25 x 0.120 = 3

mmoles of acid = H+ = 20 x 0.123 = 2.46

B +    H+ --------------> BH+

3       2.46                     0

0.54    0                         2.46

pH = pKa + log [B / BH+]

pH = 5.23 + log (0.54 / 2.46 )

pH = 4.57

D. after addition of.25.0 mL of HBr

mmoles of base pyridine = 25 x 0.120 = 3

mmoles of acid = H+ = 25 x 0.123 = 3.075

B +    H+ --------------> BH+

3       3.075                     0

0 0.075                    3

C = BH+ concentration = 3 / (25 + 25) = 0.06 M

pH = 7 -1/2 [pKb + log C]

     = 7 - 1/2 [8.77 + log 0.06]

     = 3.23

pH = 3.23

E. after addition of 32.0 mL of HBr

pH = 1.78